The record distance in the sport of throwing cowpats is 81.1 m. This record toss was set by Steve Urner of the United States in 1981. Assuming the initial launch angle was 45° and neglecting air resistance, answer the following.

(a) Determine the initial speed of the projectile.

Since only the x position and launch angle are given, I cannot figure out how to solve for initial velocity with the equations I have from the text.
Answer #1

Now the record distance is R=81.1 m Let the initial velocity of the projectile be vo The angle of projection given is 0-45 Th
Answer #2
R = 2 V^2 sinø cosø /g

V = v R . g / 2 sin45 . cos 45 = v 81.1 . 9.8 = 28.2 m/s

b) t = d / v = 81.1 / 28.2 . cos 45 = 4.068 = 4.07 s

c) The range would get less. The starting speed would be about the same. The time in the air would increase but with less horizontal speed the range would decrease